OpenStudy (anonymous):

Dan's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Dan $4.45 per pound, and type B coffee costs $5.70 per pound. This month, Dan made 123 pounds of the blend, for a total cost of $638.60 . How many pounds of type A coffee did he use?

OpenStudy (anonymous):

okay imma try to help your gonna create some equations

OpenStudy (anonymous):

a + b = 163

OpenStudy (anonymous):

and 4.45a + 5.70b = 638.60

OpenStudy (anonymous):

now rearrange the first equation I gave you to isolate for a and subsitiute into the second equation

OpenStudy (anonymous):

I do not understand @Maisa101

OpenStudy (anonymous):

do you understand the two eaquations and how i got them?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i need help solving it please and it is not 163 its 123 @Maisa101

OpenStudy (anonymous):

okay yeah sorry just fix that, okay now you know a + b = 123 can you rearrange this so that the entire equation only equals a?

OpenStudy (anonymous):

a=123b Is this right

OpenStudy (anonymous):

so very close but its not 123 times b but rather 123 - b

OpenStudy (anonymous):

now in the second equstion replace where you see the letter a with (123-b)

OpenStudy (anonymous):

What is the new equation?

OpenStudy (anonymous):

4.45(123-b)+5.70b=638.60

OpenStudy (anonymous):

okay now expand the 4.45 with the (123 -b)

OpenStudy (anonymous):

(so multiply through)

OpenStudy (anonymous):

547.35(4.45b)

OpenStudy (anonymous):

once more almost, it would actually be 547.35 - 4.45b

OpenStudy (anonymous):

now the full equation would be 547.35 - 4.45b + 5.70b = 638.60

OpenStudy (anonymous):

now can you slove for b? If thats too hard its okay start by simplifying as best as you can